# HyperD: Hybrid Periodicity Decoupling Framework for Traffic Forecasting

### v1

@@ Abstract
Accurate traffic forecasting plays a vital role in intelligent transportation systems, enabling applications such as congestion control, route planning, and urban mobility optimization. However, traffic forecasting remains challenging due to two key factors: (1) complex spatial dependencies arising from dynamic interactions between road segments and traffic sensors across the network, and (2) the coexistence of multi-scale periodic patterns (e.g., daily and weekly periodic patterns driven by human routines) with irregular fluctuations caused by unpredictable events (e.g., accidents, weather, or construction). To tackle these challenges, we propose HyperD (Hybrid Periodic Decoupling), a novel framework that decouples traffic data into periodic and residual components. The periodic component is handled by the Hybrid Periodic Representation Module, which extracts fine-grained daily and weekly patterns using learnable periodic embeddings and spatial-temporal attention. The residual component, which captures non-periodic, high-frequency fluctuations, is modeled by the Frequency-Aware Residual Representation Module, leveraging complex-valued MLP in frequency domain. To enforce semantic separation between the two components, we further introduce a Dual-View Alignment Loss, which aligns low-frequency information with the periodic branch and high-frequency information with the residual branch. Extensive experiments on four real-world traffic datasets demonstrate that HyperD achieves state-of-the-art prediction accuracy, while offering superior robustness under disturbances and improved computational efficiency compared to existing methods.
@@ Introduction
The rapid development of urbanization and sensing technologies drives abundant spatial-temporal data across domains such as traffic systems (liu2024spatial1; zhou2020foresee; ma2023rethinking), power grids (li2025causal), and environmental monitoring (liu2025spatiotemporal). Among these, traffic forecasting is a core task in intelligent transportation systems, as accurate flow prediction is essential for understanding and managing urban mobility dynamics (xia2025fairtp; zhou2025coms2t). To support such applications, forecasting models must be capable of representing rich spatial dependencies and diverse temporal patterns inherent in real-world road networks.
@@
Driven by consistent human routines and recurring activity patterns, periodicity plays a fundamental role in the temporal pattern  (zhang2017deep; yu2019citywide). As illustrated in Figure 1, traffic flow curves often repeat in consistent patterns across days and weeks, while also showing notable variation across different sensor locations. This indicates that periodicity is not only prominent but also spatially diverse, making it a critical predictive signal. As a result, many existing methods attempt to leverage these patterns through trend-seasonality decomposition  (zhou2023maintaining; yuandiffusion; wang2024towards; cao2025spatiotemporal), which separates the data into a trend component and a seasonal component. These methods suffer from two key limitations: ➀ Implicit Modeling of Periodicity — Periodic information is split across trend and seasonal components, overlooking its unified structure across time scales and spatial locations. This fragmentation hinders direct modeling of coherent periodic patterns. ➁ Inaccurate Decomposition — Traditional decomposition methods are non-learnable and inefficient, often resulting in misalignment between trend and seasonal components, and poor performance in capturing complex temporal patterns (yu2024revitalizing; kim2025comprehensive).
@@
To move beyond implicit periodic modeling, recent methods like CycleNet (cyclenet) make initial strides toward explicit periodic modeling. However, these approaches can hardly tackle traffic forecasting because they fail to model interactions between spatial locations and temporal dynamics. Additionally, they consider only a single periodic scale, leaving rich multi-scale periodic patterns underutilized. Addressing these limitations requires a flexible framework that can capture both multi-scale periodic patterns and fine-grained spatial-temporal interactions. However, periodic patterns in traffic data are often overlapping and spatially heterogeneous, while irregular fluctuations challenge conventional modeling approaches.
@@
To cope with limitations above, we propose HyperD, a Hybrid Periodic Decoupling framework that decouples traffic data into periodic and residual components, each modeled by tailored mechanisms. (1) For Implicit Modeling of periodicity, the Hybrid Periodic Representation Module uses multiple learnable embeddings and a spatial-temporal attention mechanism to model multi-scale periodic patterns. The learnable embeddings are leveraged according to the timestamps of traffic states and maintain multi-scale periodic patterns explicitly, addressing the challenge of overlapping periodic patterns that existing methods often fail to resolve. (2) For Inaccurate Decomposition, HyperD uses the Frequency-Aware Residual Representation Module to model residual dynamics, and further introduce the Dual-View Alignment Loss to ensure semantic separation between the two branches. Specifically, this loss aligns low-frequency information with the periodic branch and high-frequency information with the residual branch, encouraging each component to focus on its respective frequency band for more precise decoupling. This promotes divergence between periodic and residual components, enabling HyperD to capture both multi-scale periodic patterns and irregular fluctuations. Our main contributions are as follows:
@@
Spatial-temporal graph neural networks (STGNNs) (miao2024less; liu2024mvcar; wang2024unifying; DBLP:journals/corr/abs-2401-10134) have become a dominant approach in traffic forecasting, as they jointly model spatial correlations and temporal dynamics. The integration of LLMs with spatial-temporal models has also shown strong potential for multivariate time-series forecasting (timecma2025liu; zhang2023promptst; liu2025towards; shen2025understanding). For spatial modeling, recent approaches typically employ graph neural networks (wang2025adagcl+; GRASS; FuDHL0C25; shen2024graph; he2025mamba), to encode spatial correlations among sensors. Representative examples include STGCN (yu2018spatio) and DCRNN (li2018diffusion), which build on fixed road network structures, while GWNet (wu2019graph) and AGCRN (bai2020adaptive) further introduce adaptive graphs to capture time-varying spatial dependencies.
For temporal modeling, various architectures have been explored. RNN-based models (jiang2023spatio) capture short-term sequential patterns, while TCNs (wu2019graph; fang2021spatial) employ dilated convolutions to model longer temporal dependencies. More recently, Transformer-based models (jiang2023pdformer; gao2024spatial) leverage self-attention to flexibly capture complex temporal dynamics. However, these methods often overlook the explicit periodic patterns inherent in traffic data.
@@
To handle the complexity of traffic dynamics, some recent methods attempt to decouple spatial-temporal data into interpretable components. D2STGNN (shao2022decoupled) decouples traffic signals into diffusion and inherent components to model different propagation patterns. STWave (fang2023spatio) disentangles representations of traffic time series, separating complex traffic dynamics into stable trends and fluctuating events. STDN (cao2025spatiotemporal) decomposes traffic flow into the trend-cyclical and seasonal components in view of spatial-temporal embeddings. However, these methods capture only implicit periodic patterns.
@@
The overall framework of HyperD is shown in Figure 2. To effectively capture multi-scale periodic patterns in traffic data, we explicitly model periodicity using learnable embeddings. This allows HyperD to leverage prior knowledge of daily and weekly temporal structures, which are among the most dominant patterns in traffic data. This approach can also be easily extended to other temporal patterns, such as hourly or monthly, depending on the data. The Hybrid Periodic Representation Module consists of three main components: Learnable Daily Embeddings (LDE) and Learnable Weekly Embeddings (LWE), Spatial-Temporal Attentive Encoder (STAE), and Hybrid Periodic Pattern.
@@
To further capture long-range dependencies within the periodic patterns, we apply the self-attention mechanism along both temporal and spatial dimensions. The attention operation shares a unified form:
@@
We first project the residual component $\mathbf{R}^{\mathrm{in}}\in\mathbb{R}^{T_{1}\times N}$ into a high-dimensional space, resulting in $\tilde{\mathbf{R}}^{\mathrm{in}}\in\mathbb{R}^{T_{1}\times N\times D}$, where $D$ is the embedding dimension. To enable frequency-domain modeling, we apply Fast Fourier Transform (FFT) along both dimensions. A detailed introduction to FFT is provided in Appendix B.
@@
where $\mathcal{R}_{f}^{s}\in\mathbb{C}^{T_{1}\times N^{\prime}\times D}$ is the spatial frequency representation, and $N^{\prime}=N//2+1$ corresponds to the truncated spectrum due to symmetry in FFT.
Motivated by recent advances in frequency modeling (yi2023frequency), we introduce a complex-valued MLP (C-MLP) to refine the frequency representation. Operating in the frequency domain allows for more efficient and global modeling of non-periodic component, while reducing computational cost by focusing on essential frequency components.
@@
After obtaining the final prediction $\hat{\mathbf{Y}}$ by combining the hybrid periodic component $\mathbf{S}^{\mathrm{out}}$ and the residual component $\mathbf{R}^{\mathrm{out}}$, we need to ensure that the two components remain semantically distinct to maximize their complementary information. To achieve this, we introduce the Dual-View Alignment Loss (DVA).
@@
Datasets
To assess the performance of HyperD, we conduct experiments on four commonly used real-world traffic flow datasets: PEMS03/04/07/08 (song2020spatial). The basic statistics for these datasets are shown in Appendix C.1.
Baselines
To comprehensively evaluate our HyperD, we compare with two lines of state-of-the-art methods, including (a) spatial-temporal prediction methods: STGCN (yu2018spatio), DCRNN (li2018diffusion), GWNet (wu2019graph), ASTGCN (guo2019attention), MTGNN (wu2020connecting), STGODE (fang2021spatial), ST-WA (cirstea2022towards), DGCRN (li2023dynamic) and STPGNN (kong2024spatio), and (b) spatial-temporal decoupling methods: D2STGNN (shao2022decoupled), STWave (fang2023spatio), CycleNet (cyclenet), and STDN (cao2025spatiotemporal). We compare two CycleNet variants: CycleNet-W and CycleNet-D, both built on an MLP backbone. Detailed descriptions are provided in Appendix C.2.
Setup
To ensure a fair comparison, we adopt the experimental settings commonly used in previous studies (shao2024exploring). The datasets are divided into training, validation, and test sets with a ratio of 6:2:2, respectively. We use the past 12 time steps (previous hour) to forecast the next 12 time steps (upcoming hour). Additional implementation details are provided in Appendix C.3.
@@ Conclusion
Acknowledgemnt
This work was supported by a grant from the National Natural Science Foundation of China under grants (No.62372211, 62272191), and the Science and Technology Development Program of Jilin Province (No.20250102216JC).

### v2

@@ Abstract
Accurate traffic forecasting plays a vital role in intelligent transportation systems, enabling applications such as congestion control, route planning, and urban mobility optimization. However, traffic forecasting remains challenging due to two key factors: (1) complex spatial dependencies arising from dynamic interactions between road segments and traffic sensors across the network, and (2) the coexistence of multi-scale periodic patterns (e.g., daily and weekly periodic patterns driven by human routines) with irregular fluctuations caused by unpredictable events (e.g., accidents or weather disruptions). To tackle these challenges, we propose HyperD (Hybrid Periodic Decoupling), a novel framework that decouples traffic data into periodic and residual components. The periodic component is handled by the Hybrid Periodic Representation Module, which extracts fine-grained daily and weekly patterns using learnable periodic embeddings and spatial-temporal attention. The residual component, which captures non-periodic, high-frequency fluctuations, is modeled by the Frequency-Aware Residual Representation Module, leveraging complex-valued MLP in frequency domain. To enforce semantic separation between the two components, we further introduce a Dual-View Alignment Loss, which aligns low-frequency information with the periodic branch and high-frequency information with the residual branch. Extensive experiments on four real-world traffic datasets demonstrate that HyperD achieves state-of-the-art prediction accuracy, while offering superior robustness under disturbances and improved computational efficiency compared to existing methods.
@@ Introduction
The rapid development of urbanization and sensing technologies drives abundant spatial-temporal data across domains such as traffic systems (liu2024spatial1; zhou2020foresee; ma2023rethinking), power grids (li2025causal), and environmental monitoring (liu2025spatiotemporal). Among these, traffic forecasting is a core task in intelligent transportation systems, as accurate flow prediction is essential for understanding and managing urban mobility dynamics (xia2025fairtp; zhou2025coms2t). It requires traffic forecasting models to well capture rich spatial dependencies and diverse temporal patterns inherent in real-world road networks.
@@
Driven by consistent human routines and recurring activity patterns, periodicity plays a fundamental role in the temporal pattern of traffic data  (zhang2017deep; yu2019citywide). As illustrated in Figure 1, traffic flow curves often repeat in consistent patterns across days and weeks, while also showing notable variation across different sensor locations. This indicates that periodicity is not only prominent but also spatially diverse, making it a critical predictive signal. As a result, many existing methods attempt to leverage these patterns through trend-seasonality decomposition  (yuandiffusion; wang2024towards; cao2025spatiotemporal), which separates the data into a trend component and a seasonal component. However, these methods suffer from two key limitations: ➀ Implicit Modeling of Periodicity — Periodic information is split across trend and seasonal components, overlooking its unified structure across time scales and spatial locations. This fragmentation hinders direct modeling of coherent periodic patterns. ➁ Inaccurate Decomposition — Traditional decomposition methods are non-learnable and inefficient, often resulting in misalignment between trend and seasonal components, and poor performance in capturing complex temporal patterns (yu2024revitalizing; kim2025comprehensive).
@@
To move beyond implicit periodic modeling, recent methods like CycleNet (cyclenet) make initial strides toward explicit periodic modeling. However, these approaches can hardly tackle traffic forecasting because they fail to model interactions between spatial locations and temporal dynamics. Additionally, they consider only a single periodic scale, leaving rich multi-scale periodic patterns underutilized. Addressing these limitations requires a flexible framework that can capture both multi-scale periodic patterns and fine-grained spatial-temporal interactions. Furthermore, periodic patterns in traffic data are often overlapping and spatially heterogeneous, while irregular fluctuations challenge conventional modeling approaches.
@@
To cope with limitations above, we propose HyperD, a Hybrid Periodic Decoupling framework that decouples traffic data into periodic and residual components, each modeled by tailored mechanisms. (1) For implicit modeling of periodicity, the Hybrid Periodic Representation Module uses multiple learnable embeddings and a spatial-temporal attention mechanism to model multi-scale periodic patterns. The learnable embeddings are leveraged according to the timestamps of traffic states and maintain multi-scale periodic patterns explicitly, addressing the challenge of overlapping periodic patterns that existing methods often fail to resolve. (2) For inaccurate decomposition, HyperD uses the Frequency-Aware Residual Representation Module to model residual dynamics, and further introduce the Dual-View Alignment Loss to ensure separation between the two branches. Specifically, this loss aligns low-frequency information with the periodic branch and high-frequency information with the residual branch, encouraging each component to focus on its respective frequency band for more precise decoupling. This promotes divergence between periodic and residual components, enabling HyperD to capture both multi-scale periodic patterns and irregular fluctuations. Our main contributions are as follows:
@@
Spatial-temporal graph neural networks (STGNNs) (zhang2023mlpst; miao2024less; liu2024mvcar; wang2024unifying) have become a dominant approach in traffic forecasting, as they jointly model spatial correlations and temporal dynamics. The integration of LLMs with spatial-temporal models has also shown strong potential for multivariate time-series forecasting (timecma2025liu; zhang2023promptst; liu2025towards; shen2025understanding). For spatial modeling, recent approaches typically employ graph neural networks (wang2025adagcl+; GRASS; FuDHL0C25; shen2024graph; he2025mamba; wang2024goodat; shao2025enhanced), to encode spatial correlations among sensors. Representative examples include STGCN (yu2018spatio) and DCRNN (li2018diffusion), which build on fixed road network structures, while GWNet (wu2019graph) and AGCRN (bai2020adaptive) further introduce adaptive graphs to capture time-varying spatial dependencies.
For temporal modeling, various architectures have been explored. RNN-based models (jiang2023spatio) capture short-term sequential patterns, while TCNs (wu2019graph; fang2021spatial) employ dilated convolutions to model longer temporal dependencies. More recently, Transformer-based models (jiang2023pdformer; gao2024spatial) use attention mechanisms to capture complex temporal dynamics.
@@
To handle the complexity of traffic dynamics, some recent methods attempt to decouple spatial-temporal data into interpretable components. D2STGNN (shao2022decoupled) decouples traffic signals into diffusion and inherent components to model different propagation patterns. STWave (fang2023spatio) disentangles representations of traffic time series, separating complex traffic dynamics into stable trends and fluctuating events. STDN (cao2025spatiotemporal) decomposes traffic flow into the trend-cyclical and seasonal components in view of spatial-temporal embeddings.
@@
The overall framework of HyperD is shown in Figure 2. To effectively capture multi-scale periodic patterns in traffic data, we explicitly model periodicity using learnable embeddings. This allows HyperD to leverage prior knowledge of daily and weekly temporal structures, which are among the most dominant patterns in traffic data. This approach can also be easily extended to other temporal patterns, such as hourly or monthly. The Hybrid Periodic Representation Module consists of three main components: Learnable Daily Embeddings (LDE) and Learnable Weekly Embeddings (LWE), Spatial-Temporal Attentive Encoder (STAE), and Hybrid Periodic Pattern.
@@
To further capture long-range dependencies within the periodic patterns, we apply the self-attention mechanism (vaswani2017attention) along both temporal and spatial dimensions:
@@
We first project the residual component $\mathbf{R}^{\mathrm{in}}\in\mathbb{R}^{T_{1}\times N}$ into a high-dimensional space, resulting in $\tilde{\mathbf{R}}^{\mathrm{in}}\in\mathbb{R}^{T_{1}\times N\times D}$, where $D$ is the embedding dimension. To enable frequency-domain modeling, we apply Fast Fourier Transform (FFT) along both dimensions:
@@
where $\mathcal{R}_{f}^{s}\in\mathbb{C}^{T_{1}\times N^{\prime}\times D}$ is the spatial frequency representation, and $N^{\prime}=\left\lfloor\frac{N}{2}\right\rfloor+1$ corresponds to the truncated spectrum due to symmetry in FFT.
Motivated by recent advances in frequency modeling (yi2023frequency), we introduce a complex-valued MLP (C-MLP) to refine the frequency representation. Operating in the frequency domain with C-MLP is equivalent to convolution in the time domain, enabling more efficient and global modeling while reducing computational cost by focusing on the essential frequency components:
@@
After obtaining the final prediction $\hat{\mathbf{Y}}$ by combining the hybrid periodic component $\mathbf{S}^{\mathrm{out}}$ and the residual component $\mathbf{R}^{\mathrm{out}}$, we need to ensure the two components remain distinct to maximize their complementary information. To achieve this, we introduce the Dual-View Alignment Loss (DVA).
@@
Datasets
To assess the performance of HyperD, we conduct experiments on four commonly used real-world traffic flow datasets: PEMS03/04/07/08 (song2020spatial).
Baselines
To comprehensively evaluate our HyperD, we compare with two lines of state-of-the-art methods, including (a) spatial-temporal prediction methods: STGCN (yu2018spatio), DCRNN (li2018diffusion), GWNet (wu2019graph), ASTGCN (guo2019attention), MTGNN (wu2020connecting), STGODE (fang2021spatial), ST-WA (cirstea2022towards), DGCRN (li2023dynamic) and STPGNN (kong2024spatio), and (b) spatial-temporal decoupling methods: D2STGNN (shao2022decoupled), STWave (fang2023spatio), CycleNet (cyclenet), and STDN (cao2025spatiotemporal). We compare two CycleNet variants: CycleNet-W and CycleNet-D, both built on an MLP backbone.
@@
To ensure a fair comparison, we adopt the experimental settings commonly used in previous studies (shao2024exploring). The datasets are divided into training, validation, and test sets with a ratio of 6:2:2, respectively. We use the past 12 time steps (previous hour) to forecast the next 12 time steps (upcoming hour).
@@ Conclusion
Acknowledgment
This work was supported by a grant from the National Natural Science Foundation of China under grants (No. 62372211, 62272191), the Science and Technology Development Program of Jilin Province (No. 20250102216JC), and the China Postdoctoral Science Foundation under Grant Number (2025M771587).